OpenStudy (anonymous):

If customer A has an appointment 2 2/3 hours before customer B who has an appointment 5/6 of an hour before customer C. How much later is customer C‘s appointment than customer A’s?

OpenStudy (anonymous):

Simplify the sum of the given numbers:\[2\frac{2}{3}+\frac{5}{6}=2 + \frac{2}{3} + \frac{5}{6} = \cdots\]